A small pump takes 14 hrs to empty the storage tank. A medium pump would take 8 hrs. A large pump would take 6 hours.
How long would it take the small and medium together?

Answer by ptaylor(1894)


Let x=length of time it takes the small and medium pump working together
The small pump empties the tank at the rate of 1/14 tank per hour
The medium pump empties the tank at the rate of 1/8 tank per hour
Together they empty the tank at the rate of (1/14+1/8 or 4/56+7/56=11/56) tank per hour.
So, the question is, if both tanks working together can empty 11/56 tank per hour, how long will it take them to empty 56/56 or all of the tank?
(11/56)x=56/56 multiply each side by 56

11x=56 divide both sides by 11
x=5.090909--- hours
Hope this helps----ptaylor